Page tree

Versions Compared

Key

  • This line was added.
  • This line was removed.
  • Formatting was changed.


IssueTaskResponsibilityDue datenew date
 Completion dateComments
Modules




 

Glueing and wire bonding of 3 new IZM modulesHH,MN22.06.17


Modules should arrive end of this week/beginning of next week

6-7 Modules from PSIHH,MN29.06.17


Modules should arrive in End of next week, beginning of the week after next

Note: New modules from PSI not before June

Modules from IZM should be tested asap to order next batch (if this ones are good)

Temperature




 

Mask out of Temperatures must be possibleAA, PG05.05.1701.06.17
 

Meeting with FEB scheduled for 05.05.17 at 10am

Firmware ready, interface in development and test



MT



07.06: Screws of the external housing are missing

Repair backplanesAA,PG




MID




 

2 Micro Controllers neededAA,AK09.05.1722.05.17
 

Topic on FEB meeting on 05.05.17

Micro controllers provided but dont work as expected (see below)


MCs dont initialte M4 and M8
08.06.17


Problem most likely due to short on Master FPGA, Igor invited for 1pm.
MechanicsShort in HV feed throughAK,AD,HS,MT11.05.1718.05.17
 

Replace when first of two repaired flanges from MDC has returned (shipped already) → 23.05., Annette: 1 flange is leak tested and available for mounting; the other one has gone back today due to transportation damage

possible fixes like Y-cable proposed.


Water chiller for MID radiatorsMT



Note: only DI water

Closing of external housing



 

Missing:

Wing 2: exchange of power flange

Micro controller fix (M4,M8 issue)

Wing1:

Master FPGA

Cables test

Micro controller


Collision Switches against digonal collision Q2/Q4 AD, MT asap 
 

23.05., Annette: Draft design done, needs discussion

19.06.: 2x "Tragewinkel" ordered via ZM2, 2x "Haltewinkel" ordered via Max (MT), screws ordered via MT

21.06.: Tragewinkel announced to be ready 03.07.

comment Annette (please delete in case information is in other place or work already done):

  • do we have collision switches? (2 needed for MID, 2 needed for SPB, in case of new order: DSSC needs 4 pcs)
  • do we have vacuum compatible cable for the new switches?
  • cable on switch must be exchanged!
  • cable must be soldered on connector for Q2 and Q4 (which pin positions? → documentation)
  • communication pin positions to XFEL
  • communication logic to XFEL
 Electronics Wing2,M4 does not get signalsAK,HS,MT tbd 
  Repair after calibration of first 16 AGIPD 1.1 modules is finished, Wing 1 is tested and a bug list was created and agreed
     
  
     
  
Other




 
Calibrationimplementation of three thresholds into FPGAAA,JP,IS11.05.201701.06.17
 Separate meeting with IS needed






 








Finished tasks







Wire bonding of 5 PSI Modules and testingMN,AK05.05.1711.05.17
11.05.17

2 modules wire bonded and tests "good"

Now 4 "golden" modules available


Glue and wire bond 2 IZM modulesMN,HH11.05.17

11.05.17
  Send probed Wafers to IZM UT 18.05.17 
  Critical, one sent on 17.05.2017
MechanicsNew end switches into SPB systemMT,HS08.05.2017

 08.05.17

swap two FEMs to test 200°C issueMN, MT, HS08.05.2017

 08.05.17
Boards8 FPGA boards neededAA,AK09.05.17

 12.05.17Need Igor to return from Hospital
CablesCheck if all ordered cables were deliveredAA05.05.17

 12.05.17

Order Cable shells to fix cables to R+K profilesMT,UT11.05.1718.05.17
 Article number provided by AA on 02.05.17

Mirrored feed 8w8 ASIC power feed throughsAK,UT11.05.1718.05.17
 Might be done already by HS. Please check with him
TemperatureList of which modules show odd behavior (Jumps, xtalk from Q4 to Q3)AK,UT05.05.17

 12.05.17important to know before talking to FEB

Fix temperature issues of MID



 12.05.17

R+K profils to carry weight of new cables for full system operationMT,UT11.05.1718.05.17
 
  Repair X-ray tube HH;MT 18.05.17 
  
ElectronicsNew MFPGA firmware (delays) neededAK,UT, IS11.05.1724.05.17
 

Ongoing.

M3: One noisy ADC channel, not FIrmware related, shows squarepattern but is noisy.

M4: Signal is not received by ASICs, Possibly hardware related

M8: 8 noisy ADC channels. ADCs do not fit into range of delays (more detailed explaination needed)

Next step: swap electronics and retest, e.g. M2 with M3

  Move Detector to X-ray lab    
  
  Close down- upstream flange  AD,HS,MT  
  
  Disconnect cables AA,AK  
  All cables but 10GE optical fibres are disconnected (could not find covers)
  Call DDR   
  Comes on 18.05.17 at 10am
  Insert AGIPD 1.1 Modules   
  
  Connect Pump and cooling tubes   
  
  Connect cables (R+K Provile mandatory)   
  

Send GEL-PAKs to IZMHH24.05.17

 
 Send probed Wafers to IZM UT 24.05.17 
  how many ready? Send with GEL-PAKs?
MechanicsMount interpolatorsMT,AD08.05.201718.05.17
 Parts don't fit. New parts will be ordered by AD

Close external HousingAK,AA11.05.17 24.05.17

First MC has to be returned and it has to be verified with FEB that software upgrade does not need an open external housing

Electrical test of cablesMT,AK08.05.1718.05.17
 First test done by HS but additional electrical test needed (run module). MT takes over as long as HS is off.

XFEL interventionUT,AK,HS,MT26.05.17


outcome: Backplanes faulty

































































Recent space activity

Recently Updated
typespage, comment, blogpost
max5
hideHeadingtrue
themesocial

Space contributors

Contributors
modelist
scopedescendants
limit5
showLastTimetrue
orderupdate

Recent blog posts

Blog Posts
max4
sortmodified
contentexcerpts

Favorite Pages